OpenStudy (love.kat123):

HELP! MEDAL? find the x-intercept of the parabola with vertex (1,1) and y intercept (0,-3) write your answer in this form: (x1,y1),(x2,y2) if necessary round to the nearest hundredth

OpenStudy (freckles):

you can find the equation of the parabola then find the x-intercepts

OpenStudy (freckles):

\[y=a(x-h)^2+k \] you are given (h,k) already plug that in you can solve for a by using the other point then to find the x-intercepts you set y to 0 and solve for x

OpenStudy (freckles):

@Love.Kat123 let me know if you have any questions

OpenStudy (love.kat123):

im confused by it all math isnt my strong suit

OpenStudy (freckles):

do you know what h and k are?

OpenStudy (freckles):

(h,k) is the vertex you are given h=? and k=?

OpenStudy (freckles):

you are given the vertex is (1,1) (1,1)=(h,k) so h=? and k=?

OpenStudy (love.kat123):

1?

OpenStudy (freckles):

both are 1 yes so plug those in \[y=a(x-h)^2+k \\ y=a(x-1)^2+1\]

OpenStudy (freckles):

now find a using the other point

OpenStudy (freckles):

(x,y)=(0,-3) so replace x with 0 and y with -3 and solve for a

OpenStudy (love.kat123):

ugh im so confused and frustrated i have no idea what your talking about im legit going to cry

OpenStudy (freckles):

you should see see an x and y in my equation above

OpenStudy (freckles):

I'm asking you to replace x with 0 and y with -3

OpenStudy (freckles):

then solve for a

OpenStudy (freckles):

please show me as far as you can get with my last instruction

OpenStudy (love.kat123):

i can get about as far as reading the question i should have taken those classes more seriously

OpenStudy (freckles):

\[\color{red}y=\color{green}a(\color{blue}x-1)^2+1\] do you see the x and the y there? we are trying to use x=0 and y=-3 to find a

OpenStudy (freckles):

like can you at least replace the x with 0 and the y with -3?

OpenStudy (freckles):

like that is all I'm asking you to do right now just one step ... replacing x with 0 and y with -3

OpenStudy (love.kat123):

yes... maybe ... i think

OpenStudy (freckles):

can you show me what you have please

OpenStudy (love.kat123):

wait is x1 0 and y1 -3

OpenStudy (freckles):

though that is a y-intercept not an x-intercept

OpenStudy (freckles):

Can you please replace the x in my equation with 0 and replace the y in my equation with -3

OpenStudy (freckles):

\[\color{red}y=\color{green}a(\color{blue}x-1)^2+1 \] the y is in red the x is in blue

OpenStudy (freckles):

replace that blue letter with 0 replace that red letter with -3

OpenStudy (love.kat123):

okay so -3=a(0-1)^2+1

OpenStudy (freckles):

great so notice you have enough information to solve for a

OpenStudy (love.kat123):

wait is that right

OpenStudy (love.kat123):

it doesnt say anything about a

OpenStudy (love.kat123):

oh my im going to fail D:

OpenStudy (freckles):

yes it does... you even typed a... see the a in -3=a(0-1)^2+1

OpenStudy (freckles):

first let's do the operation next to a (0-1)^2=?

OpenStudy (love.kat123):

uh.... what

OpenStudy (freckles):

if you take 1 away from 0 you have -1... (-1)^2=?

OpenStudy (love.kat123):

i think im going to have a panic attack

OpenStudy (freckles):

you are going to have a panic attack over me asking you to evaluate (-1)^2? why?

OpenStudy (freckles):

(-1)^2 is the same as saying -1 times -1

OpenStudy (freckles):

I know you can do that

OpenStudy (love.kat123):

but the just equals -1 :|

OpenStudy (freckles):

-1 times -1 isn't -1 \[-3=a(0-1)^2+1 \\ -3=a(-1)^2+1 \\ \\ -3=a(1)+1 \\ -3=a+1\]

OpenStudy (freckles):

-1 times -1 is 1

OpenStudy (freckles):

anyways solve -3=a+1 for a

OpenStudy (love.kat123):

okay um a=-2

OpenStudy (freckles):

how did you get -2?

OpenStudy (freckles):

subtract 1 on both sides gives -3-1=a+1-1 or -3-1=a

OpenStudy (freckles):

-3-1 is?

OpenStudy (love.kat123):

you said -3=a+1 -3+1=-2

OpenStudy (love.kat123):

-3-1=-4

OpenStudy (freckles):

you cannot solve for a that way

OpenStudy (freckles):

you cannot subtract 1 on one side and do the oppsite to the other other side

OpenStudy (freckles):

-3=a+1 gives us a=-3-1 not -3+1

OpenStudy (freckles):

anyways yet a=- 4 since -4+1 is -3

OpenStudy (freckles):

\[y=a(x-h)^2+k \text{ is vertex form for a parabola } \\ (h,k)=(1,1) \\ y=a(x-1)^2+1 \\ \text{ we used} (0,-3) \text{ to find } a \\ \\ a=-4 \\ y=-4(x-1)^2+1 \\ \text{ we now have everything we need to find the } x-\text{intercepts }\]

OpenStudy (freckles):

so to finally find the x-intercepts we set y to 0 and solve for x

OpenStudy (freckles):

x-intercepts lay on the x-axis this is why y is 0 there

OpenStudy (freckles):

so you already have the y-coordinates for both your answers we are looking for the x-coordinates

OpenStudy (freckles):

and again to find those replace y with 0 and solve for x

OpenStudy (love.kat123):

whats a x-axis

OpenStudy (freckles):

the horizontal axis that is also know as the line y=0

OpenStudy (freckles):